OpenStudy (anonymous):

@whpalmer4 steve bought 12 used CDs and DVDs . CDs cost $2 each and DVDs $3. He spent $31 not including tax. How many DVDs did steve buy ?

OpenStudy (anonymous):

im thinking 7 right ?

OpenStudy (nikato):

Correct

OpenStudy (anonymous):

:)

OpenStudy (whpalmer4):

C cds and D dvds C + D = 12 2C + 3D = 31 multiply first equation by -2, and add them together: -2C - 2D = -24 2C + 3D = 31 -------------- 0C + D = 7 D = 7 C + D = 12 C + 7 = 12 C = 5 Or, the "in your head" method. Assume all 12 are CDs. 12*2 = 24. That's 31-24=7 dollars short of what he actually spent. Each DVD costs 3-2 = 1 dollar more than a CD, so trading 7 CDs into DVDs bumps up the expenditure by 7 and gives the answer.

OpenStudy (anonymous):

thank you !